OpenStudy (anonymous):

find the inverse of each function h(x)=-2x-6/5

OpenStudy (amistre64):

replace x with y and solve for "y"

OpenStudy (amistre64):

ive been staring at problems to long; it getting hard to tell if the is all over the 5 or if that is simply a 6/5 on the end. If you use () as grouping to indicate what goes together, id be helpful.

OpenStudy (anonymous):

all over the 5 srry!

OpenStudy (amistre64):

y = (-2x-6)/5 x = (-2y-6)/5 ; *5 each side 5x = -2y-6 ; +6 each side 5x +6 = -2y ; /-2 each side -(5x+6)/2 = y y = (-5x-6)/2